# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

### v1

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 21.46 22.11 21.66
Bleu score std 23.81 24.77 23.93
Bleu 1gram precision mean 21.47 22.11 21.66
Bleu 1gram precision std 23.81 24.77 23.93
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 43.57 43.54 43.59
Bleu ref len std 21.17 21.10 21.08
Bleu pred len mean 43.88 43.85 43.90
Bleu pred len std 21.21 21.14 21.11
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.00 0.00 0.00
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 21.59 22.24 21.79
Chrf score std 23.89 24.85 24.02
Ter score mean 79.33 78.68 79.13
Ter score std 24.23 25.20 24.36
Chrfpp score mean 21.59 22.24 21.79
Chrfpp score std 23.89 24.85 24.02
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.06 15.94 15.43
Bleu score std 18.77 19.65 18.65
Bleu 1gram precision mean 15.07 15.95 15.44
Bleu 1gram precision std 18.77 19.66 18.65
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 42.55 42.76 42.64
Bleu ref len std 21.06 21.06 21.01
Bleu pred len mean 42.81 43.01 42.90
Bleu pred len std 21.03 21.03 20.98
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.00 0.01 0.00
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 15.22 16.10 15.59
Chrf score std 19.06 19.92 18.94
Ter score mean 85.73 84.86 85.37
Ter score std 19.01 19.92 18.90
Chrfpp score mean 15.22 16.10 15.59
Chrfpp score std 19.06 19.92 18.94
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.59 15.21 14.92
Bleu score std 18.20 18.91 18.42
Bleu 1gram precision mean 14.60 15.21 14.92
Bleu 1gram precision std 18.20 18.91 18.42
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 43.77 43.84 43.86
Bleu ref len std 21.07 21.10 21.02
Bleu pred len mean 44.09 44.16 44.18
Bleu pred len std 21.07 21.09 21.02
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 14.76 15.37 15.08
Chrf score std 18.48 19.18 18.70
Ter score mean 86.42 85.81 86.10
Ter score std 18.55 19.28 18.78
Chrfpp score mean 14.76 15.37 15.08
Chrfpp score std 18.48 19.18 18.70
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.30 18.84 18.37
Bleu score std 21.32 22.75 22.14
Bleu 1gram precision mean 17.38 18.92 18.46
Bleu 1gram precision std 21.36 22.79 22.19
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 90.80 91.27 90.87
Bleu ref len std 45.14 45.83 45.23
Bleu pred len mean 90.83 91.38 90.85
Bleu pred len std 45.03 45.77 45.10
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 17.69 19.24 18.77
Chrf score std 21.48 22.89 22.31
Ter score mean 93.87 92.19 92.70
Ter score std 24.61 26.24 25.62
Chrfpp score mean 17.69 19.24 18.77
Chrfpp score std 21.48 22.89 22.31
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.18 18.88 18.34
Bleu score std 19.81 21.88 21.12
Bleu 1gram precision mean 17.24 18.95 18.41
Bleu 1gram precision std 19.84 21.91 21.15
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 91.72 91.87 91.81
Bleu ref len std 46.73 47.06 46.83
Bleu pred len mean 91.96 92.12 91.98
Bleu pred len std 46.59 46.92 46.68
Bleu bp mean 1.00 1.00 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 17.61 19.33 18.79
Chrf score std 20.07 22.12 21.37
Ter score mean 94.01 92.13 92.71
Ter score std 22.88 25.28 24.39
Chrfpp score mean 17.61 19.33 18.79
Chrfpp score std 20.07 22.12 21.37
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 17.25 18.33 17.12
Bleu score std 20.92 21.76 20.17
Bleu 1gram precision mean 17.32 18.39 17.19
Bleu 1gram precision std 20.96 21.80 20.23
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 92.78 92.61 92.73
Bleu ref len std 46.68 46.69 46.77
Bleu pred len mean 93.00 92.83 92.89
Bleu pred len std 46.56 46.58 46.62
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 17.65 18.73 17.52
Chrf score std 21.10 21.95 20.38
Ter score mean 93.90 92.73 94.05
Ter score std 24.29 25.30 23.46
Chrfpp score mean 17.65 18.73 17.52
Chrfpp score std 21.10 21.95 20.38
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.77 16.83 16.14
Bleu score std 19.10 20.61 19.55
Bleu 1gram precision mean 15.84 16.90 16.22
Bleu 1gram precision std 19.13 20.64 19.59
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 91.71 91.64 91.62
Bleu ref len std 46.58 46.43 46.64
Bleu pred len mean 91.83 91.76 91.69
Bleu pred len std 46.37 46.23 46.42
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03
Chrf score mean 16.17 17.23 16.54
Chrf score std 19.31 20.81 19.76
Ter score mean 95.60 94.42 95.15
Ter score std 22.20 23.92 22.73
Chrfpp score mean 16.17 17.23 16.54
Chrfpp score std 19.31 20.81 19.76
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 16.59 18.18 16.54
Bleu score std 20.30 22.44 20.48
Bleu 1gram precision mean 16.67 18.26 16.62
Bleu 1gram precision std 20.32 22.47 20.50
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 94.17 93.52 94.01
Bleu ref len std 46.56 46.42 46.60
Bleu pred len mean 94.25 93.63 94.07
Bleu pred len std 46.48 46.33 46.54
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 16.99 18.57 16.92
Chrf score std 20.48 22.59 20.64
Ter score mean 94.56 92.82 94.64
Ter score std 23.46 25.92 23.68
Chrfpp score mean 16.99 18.57 16.92
Chrfpp score std 20.48 22.59 20.64
@@
Direct Translation Prompting Refinement Word Alignment
@@
Arabic 0.84 0.84 0.83
Chinese 0.78 0.81 0.79
French 0.83 0.84 0.83
Japanese 0.78 0.82 0.78
Russian 0.81 0.82 0.80
@@
Arabic 0.84 0.85 0.84
Chinese 0.83 0.85 0.84
French 0.84 0.85 0.84
Japanese 0.85 0.86 0.85
Russian 0.82 0.83 0.80
@@
Arabic 0.86 0.86 0.85
Chinese 0.87 0.88 0.88
French 0.85 0.85 0.85
Japanese 0.90 0.90 0.90
Russian 0.84 0.84 0.81
@@
Arabic 0.84 0.85 0.84
Chinese 0.85 0.86 0.85
French 0.82 0.83 0.82
Japanese 0.88 0.88 0.88
Russian 0.82 0.82 0.80
@@
Arabic 0.82 0.83 0.81
Chinese 0.83 0.85 0.84
French 0.82 0.83 0.82
Japanese 0.89 0.89 0.89
Russian 0.82 0.83 0.80
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.42 15.12 14.66
Bleu score std 15.76 16.02 14.65
Bleu 1gram precision mean 15.54 15.21 14.76
Bleu 1gram precision std 15.83 16.06 14.68
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 169.01 177.84 176.63
Bleu ref len std 65.20 64.39 63.91
Bleu pred len mean 168.72 177.87 176.63
Bleu pred len std 64.71 64.41 64.06
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.03 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.03
Chrf score mean 15.95 15.64 15.18
Chrf score std 16.00 16.32 14.93
Ter score mean 101.29 101.47 102.00
Ter score std 20.20 20.06 18.57
Chrfpp score mean 15.95 15.64 15.18
Chrfpp score std 16.00 16.32 14.93
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 15.38 16.48 15.83
Bleu score std 16.32 17.52 16.01
Bleu 1gram precision mean 15.50 16.58 15.93
Bleu 1gram precision std 16.40 17.58 16.06
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 173.42 180.00 179.57
Bleu ref len std 66.22 64.01 63.90
Bleu pred len mean 173.41 179.91 179.47
Bleu pred len std 66.22 64.04 63.99
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02
Chrf score mean 15.90 17.00 16.36
Chrf score std 16.57 17.76 16.25
Ter score mean 101.34 99.86 100.57
Ter score std 21.12 21.79 19.97
Chrfpp score mean 15.90 17.00 16.36
Chrfpp score std 16.57 17.76 16.25
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 14.37 14.90 13.95
Bleu score std 16.32 15.58 14.30
Bleu 1gram precision mean 14.42 14.95 14.00
Bleu 1gram precision std 16.32 15.59 14.32
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 174.18 181.13 181.07
Bleu ref len std 67.36 65.07 65.15
Bleu pred len mean 174.26 181.55 181.37
Bleu pred len std 67.26 65.19 65.31
Bleu bp mean 0.99 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.01 0.02
Chrf score mean 14.82 15.37 14.40
Chrf score std 16.47 15.76 14.49
Ter score mean 102.33 101.29 102.58
Ter score std 20.70 19.21 17.92
Chrfpp score mean 14.82 15.37 14.40
Chrfpp score std 16.47 15.76 14.49

### v3

@@ Appendix C Additional Experiment Results
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 21.46 22.11 21.66 22.40
Bleu score std 23.81 24.77 23.93 25.26
Bleu 1gram precision mean 21.47 22.11 21.66 22.40
Bleu 1gram precision std 23.81 24.77 23.93 25.26
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 43.57 43.54 43.59 43.57
Bleu ref len std 21.17 21.10 21.08 21.15
Bleu pred len mean 43.88 43.85 43.90 43.88
Bleu pred len std 21.21 21.14 21.11 21.18
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.00 0.00 0.00 0.00
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 21.59 22.24 21.79 22.53
Chrf score std 23.89 24.85 24.02 25.34
Ter score mean 79.33 78.68 79.13 78.39
Ter score std 24.23 25.20 24.36 25.69
Chrfpp score mean 21.59 22.24 21.79 22.53
Chrfpp score std 23.89 24.85 24.02 25.34
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 15.06 15.94 15.43 16.31
Bleu score std 18.77 19.65 18.65 20.25
Bleu 1gram precision mean 15.07 15.95 15.44 16.31
Bleu 1gram precision std 18.77 19.66 18.65 20.25
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 42.55 42.76 42.64 42.73
Bleu ref len std 21.06 21.06 21.01 20.95
Bleu pred len mean 42.81 43.01 42.90 43.00
Bleu pred len std 21.03 21.03 20.98 20.95
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.00 0.01 0.00 0.00
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 15.22 16.10 15.59 16.47
Chrf score std 19.06 19.92 18.94 20.50
Ter score mean 85.73 84.86 85.37 84.49
Ter score std 19.01 19.92 18.90 20.52
Chrfpp score mean 15.22 16.10 15.59 16.47
Chrfpp score std 19.06 19.92 18.94 20.50
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 14.59 15.21 14.92 15.67
Bleu score std 18.20 18.91 18.42 20.06
Bleu 1gram precision mean 14.60 15.21 14.92 15.68
Bleu 1gram precision std 18.20 18.91 18.42 20.06
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 43.77 43.84 43.86 43.86
Bleu ref len std 21.07 21.10 21.02 21.10
Bleu pred len mean 44.09 44.16 44.18 44.19
Bleu pred len std 21.07 21.09 21.02 21.09
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 14.76 15.37 15.08 15.84
Chrf score std 18.48 19.18 18.70 20.31
Ter score mean 86.42 85.81 86.10 85.34
Ter score std 18.55 19.28 18.78 20.42
Chrfpp score mean 14.76 15.37 15.08 15.84
Chrfpp score std 18.48 19.18 18.70 20.31
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 17.30 18.84 18.37 18.81
Bleu score std 21.32 22.75 22.14 22.76
Bleu 1gram precision mean 17.38 18.92 18.46 18.89
Bleu 1gram precision std 21.36 22.79 22.19 22.79
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 90.80 91.27 90.87 90.93
Bleu ref len std 45.14 45.83 45.23 45.22
Bleu pred len mean 90.83 91.38 90.85 91.04
Bleu pred len std 45.03 45.77 45.10 45.15
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 17.69 19.24 18.77 19.22
Chrf score std 21.48 22.89 22.31 22.91
Ter score mean 93.87 92.19 92.70 92.21
Ter score std 24.61 26.24 25.62 26.28
Chrfpp score mean 17.69 19.24 18.77 19.22
Chrfpp score std 21.48 22.89 22.31 22.91
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 17.18 18.88 18.34 18.40
Bleu score std 19.81 21.88 21.12 21.31
Bleu 1gram precision mean 17.24 18.95 18.41 18.47
Bleu 1gram precision std 19.84 21.91 21.15 21.35
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 91.72 91.87 91.81 91.78
Bleu ref len std 46.73 47.06 46.83 46.90
Bleu pred len mean 91.96 92.12 91.98 92.01
Bleu pred len std 46.59 46.92 46.68 46.78
Bleu bp mean 1.00 1.00 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 17.61 19.33 18.79 18.84
Chrf score std 20.07 22.12 21.37 21.56
Ter score mean 94.01 92.13 92.71 92.64
Ter score std 22.88 25.28 24.39 24.64
Chrfpp score mean 17.61 19.33 18.79 18.84
Chrfpp score std 20.07 22.12 21.37 21.56
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 17.25 18.33 17.12 17.81
Bleu score std 20.92 21.76 20.17 21.04
Bleu 1gram precision mean 17.32 18.39 17.19 17.88
Bleu 1gram precision std 20.96 21.80 20.23 21.07
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 92.78 92.61 92.73 92.58
Bleu ref len std 46.68 46.69 46.77 46.78
Bleu pred len mean 93.00 92.83 92.89 92.78
Bleu pred len std 46.56 46.58 46.62 46.65
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 17.65 18.73 17.52 18.21
Chrf score std 21.10 21.95 20.38 21.23
Ter score mean 93.90 92.73 94.05 93.32
Ter score std 24.29 25.30 23.46 24.46
Chrfpp score mean 17.65 18.73 17.52 18.21
Chrfpp score std 21.10 21.95 20.38 21.23
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 15.77 16.83 16.14 16.61
Bleu score std 19.10 20.61 19.55 20.10
Bleu 1gram precision mean 15.84 16.90 16.22 16.69
Bleu 1gram precision std 19.13 20.64 19.59 20.13
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 91.71 91.64 91.62 91.33
Bleu ref len std 46.58 46.43 46.64 46.78
Bleu pred len mean 91.83 91.76 91.69 91.48
Bleu pred len std 46.37 46.23 46.42 46.58
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.01
Bleu ratio std 0.03 0.03 0.03 0.03
Chrf score mean 16.17 17.23 16.54 17.02
Chrf score std 19.31 20.81 19.76 20.32
Ter score mean 95.60 94.42 95.15 94.62
Ter score std 22.20 23.92 22.73 23.35
Chrfpp score mean 16.17 17.23 16.54 17.02
Chrfpp score std 19.31 20.81 19.76 20.32
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 16.59 18.18 16.54 17.52
Bleu score std 20.30 22.44 20.48 21.76
Bleu 1gram precision mean 16.67 18.26 16.62 17.58
Bleu 1gram precision std 20.32 22.47 20.50 21.79
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 94.17 93.52 94.01 93.31
Bleu ref len std 46.56 46.42 46.60 46.44
Bleu pred len mean 94.25 93.63 94.07 93.50
Bleu pred len std 46.48 46.33 46.54 46.36
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02 0.02
Chrf score mean 16.99 18.57 16.92 17.91
Chrf score std 20.48 22.59 20.64 21.93
Ter score mean 94.56 92.82 94.64 93.60
Ter score std 23.46 25.92 23.68 25.16
Chrfpp score mean 16.99 18.57 16.92 17.91
Chrfpp score std 20.48 22.59 20.64 21.93
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
@@
Arabic 0.84 0.84 0.83 0.85
Chinese 0.78 0.81 0.79 0.82
French 0.83 0.84 0.83 0.84
Japanese 0.78 0.82 0.78 0.83
Russian 0.81 0.82 0.80 0.83
@@
Arabic 0.84 0.85 0.84 0.85
Chinese 0.83 0.85 0.84 0.86
French 0.84 0.85 0.84 0.85
Japanese 0.85 0.86 0.85 0.87
Russian 0.82 0.83 0.80 0.84
@@
Arabic 0.86 0.86 0.85 0.86
Chinese 0.87 0.88 0.88 0.88
French 0.85 0.85 0.85 0.85
Japanese 0.90 0.90 0.90 0.90
Russian 0.84 0.84 0.81 0.84
@@
Arabic 0.84 0.85 0.84 0.85
Chinese 0.85 0.86 0.85 0.86
French 0.82 0.83 0.82 0.84
Japanese 0.88 0.88 0.88 0.89
Russian 0.82 0.82 0.80 0.83
@@
Arabic 0.82 0.83 0.81 0.83
Chinese 0.83 0.85 0.84 0.85
French 0.82 0.83 0.82 0.83
Japanese 0.89 0.89 0.89 0.89
Russian 0.82 0.83 0.80 0.83
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 15.42 15.12 14.66 15.71
Bleu score std 15.76 16.02 14.65 16.23
Bleu 1gram precision mean 15.54 15.21 14.76 15.79
Bleu 1gram precision std 15.83 16.06 14.68 16.27
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 169.01 177.84 176.63 176.93
Bleu ref len std 65.20 64.39 63.91 64.27
Bleu pred len mean 168.72 177.87 176.63 177.59
Bleu pred len std 64.71 64.41 64.06 64.61
Bleu bp mean 0.99 0.99 0.99 0.99
Bleu bp std 0.03 0.02 0.02 0.02
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.04 0.03 0.03 0.03
Chrf score mean 15.95 15.64 15.18 16.25
Chrf score std 16.00 16.32 14.93 16.54
Ter score mean 101.29 101.47 102.00 100.82
Ter score std 20.20 20.06 18.57 20.58
Chrfpp score mean 15.95 15.64 15.18 16.25
Chrfpp score std 16.00 16.32 14.93 16.54
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 15.38 16.48 15.83 17.04
Bleu score std 16.32 17.52 16.01 17.91
Bleu 1gram precision mean 15.50 16.58 15.93 17.11
Bleu 1gram precision std 16.40 17.58 16.06 17.96
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 173.42 180.00 179.57 179.61
Bleu ref len std 66.22 64.01 63.90 63.93
Bleu pred len mean 173.41 179.91 179.47 180.17
Bleu pred len std 66.22 64.04 63.99 64.31
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.03 0.02 0.02 0.02
Chrf score mean 15.90 17.00 16.36 17.58
Chrf score std 16.57 17.76 16.25 18.14
Ter score mean 101.34 99.86 100.57 99.15
Ter score std 21.12 21.79 19.97 22.40
Chrfpp score mean 15.90 17.00 16.36 17.58
Chrfpp score std 16.57 17.76 16.25 18.14
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 14.37 14.90 13.95 14.98
Bleu score std 16.32 15.58 14.30 15.59
Bleu 1gram precision mean 14.42 14.95 14.00 15.03
Bleu 1gram precision std 16.32 15.59 14.32 15.61
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 174.18 181.13 181.07 180.95
Bleu ref len std 67.36 65.07 65.15 64.85
Bleu pred len mean 174.26 181.55 181.37 181.40
Bleu pred len std 67.26 65.19 65.31 65.04
Bleu bp mean 0.99 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.01 0.02 0.01
Chrf score mean 14.82 15.37 14.40 15.45
Chrf score std 16.47 15.76 14.49 15.77
Ter score mean 102.33 101.29 102.58 101.17
Ter score std 20.70 19.21 17.92 19.26
Chrfpp score mean 14.82 15.37 14.40 15.45
Chrfpp score std 16.47 15.76 14.49 15.77